Block and Spring irn SHM with damping Due this Friday, Apr 20 at 11:59 pm (EDT) A block of mass 29.00 g on the end of spring undergoes simple harmonic motion with a frequency of 4.00 HZ What is the spring constant of the spring? 11.7 N/m Submit AnswerIncorrect. Tries 1/14 Previous Tries If the motion of the mass has an initial amplitude of 3.00 cm what is its maximum speed?. Submit Answer Tries 0/14 The amplitude decreases to 0.848 cm in 4.50 s, what is the damping constant for the system? Submit Answer Tries 0/14
Answer #1

m=29 × 10-3㎏ f 4.00Hz M2 18.3Nm b) v=2×3.14×4×0.03 v = 0.754m/ s c) 0. 00848-003e-bd%0029 Ax4%00꾜 = 0.283 bx4.50 b= 0.0163 2x